INFRARED STUDY OF THE WATER-HYDROXYL RADICAL COMPLEX TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Sensor Science Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The most prominent infrared absorptions which appear when a Ne:O_{2} mixture is codeposited at 4.3 K with a Ne:H_{2} mixture that has been passed through a microwave discharge are those of H_{2}O, HO_{2}, the HOHOH anion, and the H_{2}O(HO) complex. The absorptions of this complex correspond well with those previously obtained in argon-matrix experiments ^{a}. Photodetachment of the HOHOH anion leads to extremely great intensification of the absorptions of the H_{2}O(HO) complex. The infrared spectra of the normal and deuterium- substituted complex and the mechanisms of formation and photodestruction of the HOHOH anion will be considered.

aA. Engdahl, G. Karlström, and B. Nelander, J. Chem. Phys. 118, 7797 (2003).